OpenStudy (anonymous):

irrational numbers in algebra 2? radical 40, x^3,y

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

\[2x \sqrt{10xy}\]

OpenStudy (anonymous):

How do you get that?

OpenStudy (saifoo.khan):

\[\sqrt{40x^3y}\] \[\sqrt{4 \times 10 x^2 \times x \times y} \]

OpenStudy (saifoo.khan):

\[\sqrt{4} \sqrt{x^2} \sqrt{10xy}\] \[2 \times x \times \sqrt{10xy}\] \[2x \sqrt{10xy}\]

OpenStudy (anonymous):

\[\sqrt{20 } * \sqrt{2} * \sqrt{x^2} * \sqrt{x} * \sqrt{y}\] \[\sqrt{10} * \sqrt{2} * \sqrt{x} * \sqrt{x} * \sqrt{x} * \sqrt{y}\] \[\sqrt{5} * \sqrt{2} * \sqrt{x} * \sqrt{x} * \sqrt{x} * \sqrt{y}\] My final answer: \[5x^3y \sqrt{2}\]

OpenStudy (saifoo.khan):

INcorrect. :(

OpenStudy (anonymous):

where is my mistake?

OpenStudy (saifoo.khan):

everywhere. LOL.

OpenStudy (saifoo.khan):

http://www.twiddla.com/645968